The student government can sell a total of 400 tickets to a basketball game. An adult ticket costs $3 and a student ticket costs $2. What is the minimum number of adult tickets that must be if the student government wants to collect at least $1100?
Answer by checkley71(8403)

3x+2(400-x)=1100
3x+800-2x=1100
x=1100-800
x=300 minimum number of adult tickes needed to be sold.
proof
3*300+2*100=1100
900+200=1100.